OpenStudy (anonymous):

What is the sum of a 10–term arithmetic sequence where the first term is 3 and the last term is 75? Answer

OpenStudy (apoorvk):

sum 'S' of an AP upto 'n' where the first term 'a' and the last term 'l' is given: \[S = \frac n 2 (a+l)\]

OpenStudy (apoorvk):

**upto 'n' term^

OpenStudy (apoorvk):

plug in n=10, a =3, l=75

OpenStudy (anonymous):

ok

OpenStudy (anmolsingh):

thats easy, the Formula for sum of the terms is given by S=(n/2)(a+l) where n is the no.of terms, a is the first term and l is the last term. For your problem the n=10,a=3,l=75 That gives s=(10/2)(3+75)=390

OpenStudy (anonymous):

its 390 are you sure thats right?

OpenStudy (anmolsingh):

yes

OpenStudy (anonymous):

okay, thanks